a box has the following dimensions. what is the best value of the volume of the box? express your answers in centimetres and in inches

length= 5.21 + - 0.40 cm

width = 4.35 + - 0.30 cm

thickness = 2.30 + - 0.10 cm


ps. the symbol - (subtraction sign) is actually place down the + (addition sign)


1
Expert's answer
2021-09-21T10:56:33-0400

Assume that the best value is the greatest box. Therefore, the volume is


"V=(5.21+ 0.40)(4.35 + 0.30)(2.30 +0.10)=62.6\text{ cm}^3."

Since 1 cm is 1/2.54 of one inch, the volume in cubic inches will be


"v=\frac{62.6}{2.54^3}=3.82\text{ in}^3."